A small ferryboat is 25 m long and 10 m wide. When a container is loaded on it, the ferryboat sinks an additional 5 cm. What is the mass of the container (density of slated water = 1050 kg /m3)?
A)13,125 kg
B)36,750 kg
C)175,980 kg
D)26,250 kg
Answer #1

A) 13125 kg

let m is the mass of the container.

the weight of the displaced water when container is loaded = m*g

rho_water*V_displaced water*g = m*g

==> m = rho_water*V

= 1050*25*10*0.05

= 13125 kg